OpenStudy (anonymous):

Use the second Derivative Test to find any local extrema of the function f(x,y) =e^(4x-x^(2)-y^(2)) Please explain step by step

OpenStudy (anonymous):

\[f(x,y) = e ^{4x-x ^{2}-y ^{2}}\]

OpenStudy (anonymous):

But, you know something about the thery? Because you have to calculate the differential matrix, then check what kind of diferrential matrix you got...and so on.

OpenStudy (anonymous):

can you explain how to do the problem?

OpenStudy (amistre64):

differential matrix? .... nah

OpenStudy (amistre64):

what do you know about derivative rules?

OpenStudy (anonymous):

i got fx = e^4x-x^2-y^2(4-2x) and fy= 4x-x^2-y^2(2y)

OpenStudy (anonymous):

but i dont know what to do after that

OpenStudy (amistre64):

well, youve got good form. with the tiny little exponent i didnt see that ys :) take Fxx and Fyy and Fxy i believe to find the extrema of the surface

OpenStudy (anonymous):

I dont know how to get any of those in my book it says to set fx=0 and fy=0 and solve?

OpenStudy (amistre64):

you know how to find Fx and Fy; find the x again of Fx to get Fxx, find the y again of Fy to get Fyy

OpenStudy (anonymous):

i understand you get fxx and fyy by taking the derivative of fx and fy respectively but i keep getting the wrong answer

OpenStudy (amistre64):

the wrong derivatives, or the wrong answer?

OpenStudy (anonymous):

f(x,y) = x^2y^3 then fx = 2x.y^2 Then fxy = 2x.2y

OpenStudy (anonymous):

both

OpenStudy (amistre64):

let me try it; if anything just for the practice :)

OpenStudy (amistre64):

\[f(x,y) = e ^{4x-x ^{2}-y ^{2}}\] \[fx=(4-2x)e^{4x-x ^{2}-y ^{2}};\ fy=(-2y)e^{4x-x ^{2}-y ^{2}}\] \[fxx=(-2)e^{4x-x ^{2}-y ^{2}}+(4-2x)^2e^{4x-x ^{2}-y ^{2}}\]\[ \ fyy=(-2)e^{4x-x ^{2}-y ^{2}}+(-2y)^2e^{4x-x ^{2}-y ^{2}}\] good so far?

OpenStudy (amistre64):

now, the zeros of this would give us some critical values, but with a 3d surface we might be stuck in a saddle point and so we have a way to check for that

OpenStudy (anonymous):

Don´t forget that your critical values will be given by the gradient (fx, fy) = 0, 0

OpenStudy (amistre64):

D=Fxx*Fyy - (Fxy)^2

OpenStudy (anonymous):

i'm having trouble comming up with fxx and fyy did you distribute the 4-2x and 2y into e^4x^2x^2-2y?

OpenStudy (amistre64):

no, i just used it as a product and pulled it thru the product rule

OpenStudy (amistre64):

[fg]'=f'g+fg'

OpenStudy (anonymous):

ohallright that makes total sense x.x

OpenStudy (amistre64):

http://mysite.science.uottawa.ca/cstar050/1300notes/lecture19.pdf this might be a good review if you need it

OpenStudy (anonymous):

thank you, it seems really complicated though to multiply fxxfyy and subtract it from fxy^2

OpenStudy (amistre64):

wolframalpha.com would help to ease that burden ;)

OpenStudy (amistre64):

\[fx=(4-2x)e^{4x-x ^{2}-y ^{2}}\] since 4-2x has no y in it; its considered a constant and we get: \[fxy=(-2y)(4-2x)e^{4x-x ^{2}-y ^{2}}\] \[(fxy)^2=((-2y)(4-2x)e^{4x-x ^{2}-y ^{2}})^2\] \[(fxy)^2=(4y^2)(16-16x+4x^2)e^{(4x-x ^{2}-y ^{2})^2}\] \[fxx*fyy=(14-16x+4x^2)(-2+4y^2)e^{(4x-x ^{2}-y ^{2})^2}\] \[fxx*fyy=(14-16x+4x^2)(-2+4y^2)e^{(4x-x ^{2}-y ^{2})^2}\]\[-(fxy)^2=-(4y^2)(16-16x+4x^2)e^{(4x-x ^{2}-y ^{2})^2}\] \[[(14-16x+4x^2)(-2+4y^2)-(4y^2)(16-16x+4x^2)]e^{(4x-x ^{2}-y ^{2})^2}\] whew!! it just takes some time

OpenStudy (anonymous):

its super long =/ but doesnt plugging in 0 for that give me 14(-2)-0?

OpenStudy (amistre64):

if you plug in 0,0; yes

OpenStudy (amistre64):

how does that compare to Fxx?

OpenStudy (anonymous):

but that would mean that it is -28<0 so there is no max or min at (0,0) and the answer is e^4 =local max

OpenStudy (anonymous):

and at fxx(0,0) i get 14which is >0

OpenStudy (amistre64):

D > 0 and fxx(a,b) < 0 ; MAX D > 0 and fxx(a,b) > 0 ; MIN D<0; SADDLE

OpenStudy (anonymous):

at (1,1) im getting -12e^4 for D and for fxx 2e^2 i dont think thats right

OpenStudy (amistre64):

fx=0 when x=2 fy=0 when y=0 ; our critical point is at (2,0)

OpenStudy (anonymous):

i get -2e^4 for fxx =/

OpenStudy (amistre64):

\[fxx(2,0)=−2e^{4}\]\[fxy(2,0)=0\]

OpenStudy (amistre64):

D = (14−16(2)+4(4))(−2) D = (14−32+16)(−2) D = (-2)(−2) = 4 D>0, and Fxx < 0 this would be a max right?

OpenStudy (amistre64):

graphing it would be nice :)

OpenStudy (anonymous):

the max is e^4 isnt that supposed to be our answer?

OpenStudy (amistre64):

e^4 might be the max, but graphing it tells us that the max does indeed occur at (2,0) e^(4x-x^2-y^2); (2,1) = e^4

OpenStudy (amistre64):

(2,0) ... my keyboard hates me

OpenStudy (anonymous):

lol dw its not you it's this problem it hates xD allright thanks so much! I know it was a headache, but I appreciate the help!

OpenStudy (amistre64):

good luck with it all ;)